10. Solve the system of equations: x + 2y = 3
4x + 5y = 6

Answer by MRperkins(300)

x+2y=3
4x+5y=6
x=3-2y
4(3-2y)+5y=6
12-8y+5y=6
12-3y=6
-3y=-6
y=2
x+2(2)=3
x+4=3
x=-1
check solution x=-1, y=2 or (-1,2) written as a coordinate point.
-1+2(2)=3
3=3 so it works in this equation
4(-1)+5(2)=6
6=6 so it works in this equation. Therefore, your solution is (-1,2)
Email me at justin.sheppard.tech@hotmail.com for more help. We can even set up an online whiteboard to practice problems.
Good luck
Justin